Calculate the inbreeding coefficient (F) for a population that has allele frequencies of p=0.6 and q=0.4, and has 500 individuals that are homozygous for p, 100 individuals that are heterozygous, and 400 individuals that are homozygous for q.
Answer #1

inbreeding coefficient (F) = # of expected heterozygotes - # of observed heterozygote / # of expected heterozygotes

from Hardy-Weinberg principle - 2pq (# of heterozygotes) = 2 * 0.6 * 0.4 = 0.48

Since total number of individuals are 1000, multiply the above value by 1000 to get # of expected heterozygotes-

# of expected heterozygotes = 0.48 * 500 = 480

# of observed heterozygote = 100

put all the value in formula

F = 480-100 / 480 = 0.791